Solve by the elimination method.
3x + 4y = 3
6x + 8y = 6
Can someone please help me with this. I was doing really well with my work but now am stuck again. Help me please.

Answer by checkley77(12844)

3x+4y=3 multiply this equation by -2 & add the 2 equations.
-6x-8y=-6
6x+8y=6
-----------------------
0+0=0 These 2 equations are identical with 2 unknowns. Thus there is no unique solution